# Diffusion Models, Image Super-Resolution And Everything: A Survey

### v2

@@ V Diffusion Models for Image SR
So far, our discussion has revolved around the theoretical framework of DMs. This section explores the practical applications and recent publications in image SR. We will discuss concrete realizations of diffusion models, which are predominantly DDPMs. Next, we will introduce alternative input domains for DDPMs and strategies to incorporate various ways of conditioning information (LR image) and how to improve conditioning with guidance. Moreover, we will explore zero-shot approaches, as well as for SR specifically interesting research areas, such as corruption spaces, color-shifting and architectural designs.
@@ V-A Concrete Realization of Diffusion Models
While SGMs provide considerable flexibility in design, the prevailing trend within the field of image leans towards the implementation of DDPMs. Firstly, DDPMs are characterized by their straightforward implementation, which reduces the barrier to entry for researchers and developers alike. This ease of implementation is a significant advantage, as it allows for quicker development cycles and easier replication of results. Additionally, the inherent flexibility offered by SGMs, while advantageous in creating customized solutions, introduces complexity due to the multitude of design variables that need to be considered. This complexity can pose a challenge in research settings, where rigorously evaluating the impact of each variable (e.g., different sampling algorithms) becomes a cumbersome process. Moreover, the growing body of literature featuring DDPM implementations has contributed to their increasing popularity within the community. As more studies adopt and validate DDPMs for image SR applications, a virtuous cycle is created, where the familiarity and proven effectiveness of DDPMs encourage further adoption. This trend towards DDPMs is further reinforced by cumulative knowledge and shared experiences, making them a more common choice for image SR.
Among the pioneering DM efforts in SR is SR3 [15], a method that concretely realizes DDPMs for image SR. Like typical in DDPMs, it adds Gaussian noise to the LR image until $\mathbf{z}_{T}\sim\mathcal{N}\left(\mathbf{0},\mathbf{I}\right)$ and generates a target HR image $\mathbf{z}_{0}$ iteratively in $T$ refinement steps. SR3 employs the denoising model to predict the noise $\epsilon_{t}$. The denoising model, $\varphi_{\theta}\left(\mathbf{x},\mathbf{z}_{t},\gamma_{t}\right)$, takes the LR image $\mathbf{x}$, the noise variance $\gamma_{t}$, and the noisy target image $\mathbf{z}_{t}$ as inputs.
With the prediction of $\epsilon_{t}$ provided by $\varphi_{\theta}$, we can reformulate Equation 14 to approximate $\mathbf{z}_{0}$ as follows:
@@
Concurrent work focused on a similar implementation of SR3 but shows different variations implementing the denoising model $\varphi_{\theta}\left(\mathbf{x},\mathbf{z}_{t},\gamma_{t}\right)$, which we will discuss later.
@@
A notable mention is SRDiff [76], published around the same time and follows a close realization of SR3. The main distinction between SRDiff and SR3 is that SR3 predicts the HR image directly, whereas SRDiff predicts the residual information between the LR and HR image, i.e., the difference. Thus, it focuses on an alternative target domain, which is part of the next section.
V-B Alternative Domains
@@
So far, we have discussed methods that operate on the pixel space directly. This section will look at different methods that tackle the problem by mapping it to latent and frequency space. Apart from particular challenges arising from the alternative input domain, these methods incur an additional step that maps the pixel domain into their own, as illustrated in Figure 3. Moreover, we will introduce the residual target domain.
@@
Figure 3: Overview of alternative domains. The green bar shows the vanilla DM that operates in pixel space. The blue bar shows the exploit of the latent space domain via Autoencoders. The red bar shows the application of DMs within the wavelet domain.
V-B1 Latent Space Domain
Models like SR3 [15] and SRDiff [76] have achieved high-quality SR results by operating in the pixel domain. However, these models are computationally intensive due to their iterative nature and the high-dimensional calculations in RGB space. To reduce computational demands, Latent Diffusion Models (LDMs), introduced by Rombach et al. [13], perform diffusion processes in a low-dimensional latent space. This approach significantly lowers resource requirements without compromising performance and allows the reuse of latent representations across multiple models. LDMs typically follow a two-stage process. First, a pre-trained autoencoder, like the VQ-GAN [9], compresses input data into a low-dimensional latent representation. This encoder is trained using perceptual and adversarial losses. In the second stage, a time-conditional U-Net model is trained in the latent space, incorporating image-specific biases. For image SR, low-resolution images are concatenated as a condition.
Similarly, Latent Score-based Generative Models (LSGMs), as introduced by Vadhat et al. [77], apply SGMs within the latent space. Unlike LDMs, they jointly train their SGM with a VAE. They achieve faster sampling speeds by pre-training their VAE. Another latent model, but for SDEs, is presented by Luo et al. [78, 79] with Refusion (image Restoration with diffusion models). It employs a mean-reverting Image Restoration (IR)-SDE to transform HR images into their degraded counterparts. They leverage an autoencoder to compress the LR image into its latent representation, with skip connections for accessing multi-scale details. Also, the authors introduce nonlinear activation-free blocks (NAFBlocks) [80] into the IR-SDE [79]. These NAFBlocks replace all non-linear activations with an element-wise operation that splits feature channels into two parts and multiplies them to produce one output.
Chen et al. [81] employ a similar approach by proposing a two-stage strategy called the Hierarchical Integration Diffusion Model (HI-Diff). In the first stage, a Latent Encoder (LE) compresses the ground truth image to a highly compact latent space representation. This is treated as a prior feature to a Hierarchical Integration Module (HIM) transformer jointly trained with the LE using L1 loss. A DM is trained to generate the prior feature in the next stage.
V-B2 Wavelet Target Domain
@@
Wavelets provide a novel outlook on SR [16, 82]. The conversion from the spatial to the wavelet domain is lossless and offers significant advantages. The spatial size of an image can be downsized by a factor of four, thereby allowing faster diffusion during the training and inference stages. Moreover, wavelet transformations naturally segregate high-frequency details into distinct channels, which facilitates a more concentrated and intentional focus on high-frequency information, offering a higher degree of control during the SR process [83]. Besides, it can be conveniently incorporated into existing DMs as a plug-in feature.
The diffusion process can interact directly with all wavelet bands as proposed in DiWa [84] or specifically target certain bands while the remaining bands are predicted via standard CNNs. For instance, WaveDM [85] modifies the low-frequency band, whereas WSGM [86] or ResDiff [87] conditions the high-frequency bands relative to the low-resolution image. Altogether, the wavelet domain presents a promising avenue for future research. It provides potential for significant performance acceleration while maintaining, if not enhancing, the quality of SR results.
Figure 4: Overview of different ways to condition a diffusion-based SR model. One can incorporate the LR image, the SR image, its features, or a text-embedding of the LR image during the backward diffusion. Note that also combinations of different conditioning methods are possible.
V-B3 Residual Target Domain
As mentioned, SRDiff [76] is the pioneering work that advocates for shifting the image generation process into the residual space. This strategy aims to generate the difference between the upsampled LR and the HR image. This enables the DM to focus on residual details, speeds up convergence, and stabilizes the training [16]. Another notable contribution is ResDiff [87], which also seeks to predict the residuals. Whang et al. [88] also employs residual predictions as a fundamental component of their predict-and-refine approach for image deblurring.
In a different vein, Yue et al. [89] presents a technique that constructs a Markov chain of transformations between HR and LR images by manipulating the residual between them. This novel approach substantially enhances transition efficiency, providing an innovative perspective on image manipulation strategies.
V-C Conditioning Diffusion Models
@@
DMs heavily depend on conditioning information to guide the sampling process toward a reasonable HR prediction. One common strategy is to incorporate the LR image during the backward diffusion. However, there are also other strategies that vary from the standard protocol. This section reviews various methods (see Figure 4) for seamlessly integrating conditioning information into backward diffusion.
V-C1 Low Resolution Reference
@@
As SR3 [90] exemplifies, high-quality SR predictions can be achieved through a straightforward channel concatenation. The LR image is concatenated with the denoised result from time step $t-1$ and serves as the conditioning input for noise prediction at time step $t$. In contrast, Iterative Latent Variable Refinement (ILVR) by Choi et al. [91] conditions the generative process of an unconditional LDM. This approach offers the advantage of shorter training times, as it leverages a pre-trained DM. To integrate conditioning information, the low-frequency components of the denoised output are replaced with their corresponding counterparts from the LR image. Thus, the latent variable is aligned with a provided reference image at each generation process stage, ensuring precise control and adaptation during generation.
V-C2 Super-Resolved Reference
@@
An alternative to conditioning the denoising on the LR image involves learned priors from pre-trained SR models to predict a reference image. E.g., CDPMSR [92] conditions the denoising process with a predicted SR reference image obtained using existing and standard SR models. ResDiff [87], on the other hand, leverages a pre-trained CNN to predict a low-frequency, content-rich image that includes partial high-frequency components. This image guides the noise towards the residual space, offering an alternative means of conditioning the generative process.
@@ V-C3 Feature Reference
Another avenue for conditioning involves relevant features extracted from pre-trained networks. SRDiff [76] leverages a pre-trained encoder to encode LR image features at each step of the backward diffusion. These features serve as guidance, aiding in the generation of higher-resolution outputs. Implicit DMs (IDMs) [75] take a different approach by conditioning their denoising network with a neural representation, which enables the learning of a continuous representation at various scales. They encode the image as a function within continuous space and seamlessly integrate it into the DM. These extracted features are adapted to multiple scales and are used across multiple layers within the DMs. To comprehensively understand the performance differences between these approaches, comparisons can be found in Table II and Table I.
V-C4 Exploiting Text-To-Image
By incorporating conditioning information that goes beyond the LR image (e.g., its SR prediction, direct concatenation of the LR image, or its feature representation), one can add Text-To-Image (T2I) information. The incorporation of T2I information proves advantageous as it allows the usage of pre-trained T2I models. These models can be fine-tuned by adding specific layers or encoders tailored to the SR task, facilitating the integration of textual descriptions into the image generation process. This approach enables a richer source of guidance, potentially leading to improved image synthesis and interpretation in SR tasks.
@@ V-C4 Exploiting Text-To-Image
Wang et al. have put this concept into practice with StableSR [100]. Central to StableSR is a time-aware encoder trained in tandem with a frozen Stable DM, essentially a LDM. This setup seamlessly integrates trainable spatial feature transform layers, enabling conditioning based on the input image. To further augment the flexibility of StableSR and achieve a delicate balance between realism and fidelity, they introduce an optional controllable feature wrapping module. This module accommodates user preferences, allowing for fine-tuned adjustments based on individual requirements. The inspiration for this feature comes from the methodology introduced in CodeFormer [101], which enhances the versatility of StableSR in catering to diverse user needs and preferences.
Likewise, Yang et al. introduce a method known as Pixel-Aware Stable Diffusion (PASD) [102]. PASD takes conditioning a step further by incorporating text embeddings of the LR input using a CLIP text encoder [103] and its feature representation. This approach augments the model’s ability to generate images by incorporating textual information, thus allowing for more precise and context-aware image synthesis. Comparisons between PASD and other approaches can be found in Table III, demonstrating the impact of this text-based conditioning on image SR results.
@@
V-D Guidance in Training
We have examined the learning of conditional distributions, which form the backbone of image SR [109, 15]. As such, the condition $\mathbf{x}$, such as an LR image, is integrated into the backward diffusion, i.e., $p_{\theta}\left(\mathbf{z}_{t-1}\mid\mathbf{z}_{t},\mathbf{x}\right)$ for DDPMs or in $s_{\theta}(\mathbf{z}_{t},\mathbf{x},t)$ for SGMs/SDEs. However, this simple formulation can result in a model that overlooks the conditioning.
A principle known as guidance can mitigate this issue by controlling the weighting of the conditioning information at the expense of sample diversity. This principle is categorized into Classifier and Classifier-Free guidance. To our knowledge, while effectively used for improving DMs, they have not been applied to the task of SR. We include a brief overview of these methods as they comprise an exciting avenue for future work on image SR.
V-D1 Classifier Guidance
Classifier guidance employs a classifier to guide the diffusion process by merging the score estimate of the DM with the gradients of the classifier during sampling [12]. This process is similar to low temperature or truncated sampling in BigGANs [110] and facilitates a trade-off between mode coverage and sample fidelity. The classifier is trained concurrently with the DM to predict the conditional information $\mathbf{x}$ from $\mathbf{z}_{t}$. For weighting of the conditioning information, the score function becomes:
V-D2 Classifier-Free Guidance
Classifier-Free guidance aims to achieve similar results without a classifier [111]. It exploits the vanilla formulation such that Equation 44 can be rearranged into

### v3

@@ V Diffusion Models for Image SR
So far, we introduced the theoretical framework of DMs. This section reviews practical applications and recent advances in image SR. We will discuss concrete realizations of DMs, which are predominantly DDPMs. We then discuss guidance strategies to enhance conditioning usage, represent conditioning information in alternative state domains for DDPMs, and incorporate various conditioning methods. Additionally, we explore SR-specific research areas, including corruption spaces, color-shifting, and architectural designs. Figure 3 provides a topological overview of this section.
Figure 3: Topology of this work. Conditioning (subsection V-D) leads the backward diffusion, whereas guidance (subsection V-B) is a training strategy to improve the incorporation of conditioning into DMs. The state domain (subsection V-C) describes the representation of states $\mathbf{z}_{t}$. The corruption space (subsection V-E) describes the target of the forward diffusion process or the start of the backward diffusion.
@@ V-A Concrete Realization of Diffusion Models
While SGMs provide considerable design flexibility, the image SR trend leans towards DDPMs. DDPMs benefit from a straightforward implementation, which reduces the entry barrier. It is a significant advantage, as it allows quicker development cycles and replication of results. In addition, while the flexibility of SGMs is advantageous in creating customized solutions, it introduces design complexity due to the multitude of design variables that need to be considered. This poses a challenge in research settings, where rigorously evaluating the impact of each variable (e.g., different sampling algorithms) becomes cumbersome. Moreover, the growing DDPM literature contributes to their popularity. As more studies adopt DDPMs, a virtuous cycle is created, where familiarity and proven effectiveness encourage further adoption.
Among the pioneering DM efforts is SR3 [15], which concretely realizes DDPMs for image SR. Like typical for DDPMs, it adds Gaussian noise to the LR image until $\mathbf{z}_{T}\sim\mathcal{N}\left(\mathbf{0},\mathbf{I}\right)$ and generates a target HR image $\mathbf{z}_{0}$ iteratively in $T$ refinement steps. SR3 employs the denoising model to predict the noise $\epsilon_{t}$. The denoising model, $\varphi_{\theta}\left(\mathbf{x},\mathbf{z}_{t},\gamma_{t}\right)$, takes the LR image $\mathbf{x}$, the noise variance $\gamma_{t}$, and the noisy target image $\mathbf{z}_{t}$ as inputs. With the prediction of $\epsilon_{t}$ provided by $\varphi_{\theta}$, we can reformulate Equation 14 to approximate $\mathbf{z}_{0}$ as follows:
@@
Concurrent work focused on a similar implementation of SR3 but shows different variations implementing the denoising model $\varphi_{\theta}\left(\mathbf{x},\mathbf{z}_{t},\gamma_{t}\right)$, which we will discuss later. A notable mention is SRDiff [79], published around the same time and follows a close realization of SR3. The main distinction between SRDiff and SR3 is that SR3 predicts the HR image directly, whereas SRDiff predicts the residual information between the LR and HR image, i.e., the difference. Thus, it has an alternative state domain, which will be discussed next.
V-B Guidance in Training
The backbone of diffusion-based image SR is the learning of conditional distributions [101, 15]. As such, the condition $\mathbf{x}$, e.g., the LR image, is integrated into the backward diffusion, i.e., $p_{\theta}\left(\mathbf{z}_{t-1}\mid\mathbf{z}_{t},\mathbf{x}\right)$ for DDPMs or in $s_{\theta}(\mathbf{z}_{t},\mathbf{x},t)$ for SGMs/SDEs. However, this simple formulation can result in a model that overlooks the conditioning. A principle known as guidance can mitigate this issue by controlling the weighting of the conditioning information at the expense of sample diversity. It can be categorized into classifier and classifier-free guidance. To our knowledge, while effectively used for improving DMs, they have not been applied to image SR.
Classifier Guidance: Classifier guidance employs a classifier to guide the diffusion process by merging the score estimate of the DM with the gradients of the classifier during sampling [12]. This process is similar to low temperature or truncated sampling in BigGANs [102] and facilitates a trade-off between mode coverage and sample fidelity. The classifier is trained concurrently with the DM to predict the conditional information $\mathbf{x}$ from $\mathbf{z}_{t}$. For weighting of the conditioning information, the score function becomes:
Classifier-Free Guidance: Classifier-Free guidance aims to achieve similar results without a classifier [103]. It modifies Equation 32 into
V-C State Domains
@@
So far, we have discussed methods that operate directly on the pixel space. This section introduces different methods that map the input into alternative state domains: latent, frequency, and residual space. Apart from particular challenges arising from the alternative state domain, these methods incur an additional step that maps the pixel domain into their own, as illustrated in Figure 4.
Latent Space: Models like SR3 [15], and SRDiff [79] have achieved high-quality SR results by operating in the pixel domain. However, these models are computationally intensive due to their iterative nature and the high-dimensional calculations in RGB space. To reduce computational demands, one can move the diffusion process into the latent space of an autoencoder [105]. The first of this kind was the Latent Score-based Generative Models (LSGMs) by Vadhat et al. [106]. It is a regular SGM that operates in the latent space of a VAE and, by pre-training the VAE, achieves even faster sampling speeds. It yields comparable and better results than DMs operating in the pixel domain while being faster. Building upon LSGMs, Rombach et al. introduced the Latent Diffusion Model (LDM) [107, 13], which also performs diffusion in a low-dimensional latent space of an autoencoder. In contrast to LSGM, LDM utilizes a DDPM and an autoencoder that is pre-trained, like the VQ-GAN [5], and is not jointly trained with the denoising network. This approach significantly lowers resource requirements without compromising performance. Due to the decoupled training, it requires very little regularization of the latent space and allows the reuse of latent representations across multiple models. Improving upon LDMs is REFUSION (image REstoration with difFUSION models) [108] by Luo et al., which differs in two aspects: First, it uses a U-Net that contains skip connections from the encoder to the decoder, which provides the decoder with additional details. Moreover, it introduces Nonlinear Activation-Free blocks (NAFBlocks) [109], replacing all non-linear activations with an element-wise operation that splits feature channels into two parts and multiplies them to produce one output. Secondly, they train their U-Net with a latent-replacing training strategy, which partially replaces the latent representation with either the encoded LR or HR image for reconstruction training. Similarly, Chen et al. [110] improve the architectural aspects of LDMs and propose a two-stage strategy called the Hierarchical Integration Diffusion Model (HI-Diff). In the first stage, an encoder compresses the ground truth image to a highly compact latent space representation, which has a much higher compression ratio than LDM. As a result, the computational burden of the DM, which refines multi-scale latent representations, is much more reduced. The second stage is a vision transformer-based autoencoder, which incorporates the latent representations of the first stage during the downsampling process via Hierarchical Integration Modules (HIM), a cross-attention fusion module.
@@
Figure 4: Overview of state domains. The green bar shows the vanilla DM operating in pixel space. The blue bar shows the exploit of the latent space domain via Autoencoders. The red bar shows the application of DMs in the wavelet domain.
@@
Frequency Space: Wavelets provide a novel outlook on SR [16, 111]. The conversion from the spatial to the wavelet domain is lossless and offers significant advantages as the spatial size of an image can be downsized by a factor of four, thereby allowing faster diffusion during the training and inference stages. Moreover, the conversion segregates high-frequency details into distinct channels, facilitating a more concentrated and intentional focus on high-frequency information, offering a higher degree of control [112]. Besides, it can be conveniently incorporated into existing DMs as a plug-in feature. The diffusion process can interact directly with all wavelet bands as proposed in DiWa [113] or specifically target certain bands while the remaining bands are predicted via standard CNNs. For instance, WaveDM [114] modifies the low-frequency band, whereas WSGM [115] or ResDiff [116] conditions the high-frequency bands relative to the low-resolution image. Altogether, the wavelet domain presents a promising avenue for future research. It provides potential for significant performance acceleration while maintaining, if not enhancing, the quality of SR results.
Residual Space: SRDiff [79] was the first work that advocated for shifting the generation process into the residual space, i.e., the difference between the upsampled LR and the HR image. This enables the DM to focus on residual details, speeds up convergence, and stabilizes the training [16, 111]. Whang et al. [117] also employs residual predictions as a fundamental component of their predict-and-refine approach for image deblurring. However, unlike SRDiff, they provide a SR prediction with a CNN instead of the bilinear upsampled LR and predict the residuals between the SR prediction and the HR ground truth with their DM. An improvement is presented by ResDiff [116], which additionally incorporates the SR prediction and its high-frequency information during the backward diffusion for better guidance. In a different vein, Yue et al. [118] presents ResShift. This technique constructs a Markov chain of transformations between HR and LR images by manipulating the residual between them. Thus, instead of just adding Gaussian noise with zero mean in the forward process, the residual is also added as the mean of the noise sampling during training. This novel approach substantially enhances sampling efficiency, i.e., only 15 sampling steps.
V-D Conditioning Diffusion Models
@@
DMs depend on conditioning information to guide the sampling process toward a reasonable HR prediction. One common strategy is to use the LR image during the backward diffusion. This section reviews various alternative methods for integrating conditioning information into backward diffusion.
@@
Low Resolution Reference: High-quality SR predictions can be achieved through a straightforward channel concatenation [119]. The LR image is concatenated with the denoised result from time step $t-1$ and serves as the conditioning input for noise prediction at time step $t$. In contrast, Iterative Latent Variable Refinement (ILVR) by Choi et al. [120] conditions the generative process of an unconditional LDM [13]. This approach offers the advantage of shorter training times, as it leverages a pre-trained DM. To integrate conditioning information, the low-frequency components of the denoised output are replaced with their corresponding counterparts from the LR image. Thus, the latent variable is aligned with a provided reference image at each generation process stage, ensuring precise control and adaptation during generation.
@@
Super-Resolved Reference: An alternative to conditioning the denoising on the LR image involves learned priors from pre-trained SR models to predict a reference image. E.g., CDPMSR [121] conditions the denoising process with a predicted SR reference image obtained using existing and standard SR models. ResDiff [116], on the other hand, leverages a pre-trained CNN to predict a low-frequency, content-rich image that includes partial high-frequency components. This image guides the noise towards the residual space, offering an alternative means of conditioning the generative process.
@@
Feature Reference: Another avenue for conditioning involves relevant features extracted from pre-trained networks. SRDiff [79] leverages a pre-trained encoder to encode LR image features at each step of the backward diffusion. These features serve as guidance, aiding in the generation of higher-resolution outputs. Implicit DMs (IDMs) [100] take a different approach by conditioning their denoising network with a neural representation, which enables the learning of a continuous representation at various scales. They encode the image as a function within continuous space and seamlessly integrate it into the DM. These extracted features are adapted to multiple scales and are used across multiple layers within the DMs. To comprehensively understand the performance differences between these approaches, comparisons can be found in Table II and Table I. Recently, DeeDSR was introduced [129], which incorporates degradation-aware features extracted from the LR image to guide the diffusion process of a LDM [107].
@@
Text-to-Image Information: By incorporating conditioning information that goes beyond the LR image (e.g., its SR prediction, direct concatenation of the LR image, or its feature representation), one can add Text-To-Image (T2I) information. The incorporation of T2I information proves advantageous as it allows the usage of pre-trained T2I models. These models can be fine-tuned by adding specific layers or encoders tailored to the SR task, facilitating the integration of textual descriptions into the image generation process. This approach enables a richer source of guidance, potentially improving image synthesis and interpretation in SR tasks. Wang et al. have put this concept into practice with StableSR [107]. Central to StableSR is a time-aware encoder trained in tandem with a frozen Stable DM, essentially a LDM. This setup seamlessly integrates trainable spatial feature transform layers, enabling conditioning based on the input image. To further augment the flexibility of StableSR and achieve a delicate balance between realism and fidelity, they introduce an optional controllable feature wrapping module. This module accommodates user preferences, allowing for fine-tuned adjustments based on individual requirements. The inspiration for this feature comes from the methodology introduced in CodeFormer [130], which enhances the versatility of StableSR in catering to diverse user needs and preferences. Likewise, Yang et al. introduce a method known as Pixel-Aware Stable Diffusion (PASD) [131]. PASD takes conditioning a step further by incorporating text embeddings of the LR input using a CLIP text encoder [59] and its feature representation. This approach augments the model’s ability to generate images by incorporating textual information, thus allowing for more precise and context-aware image synthesis. Comparisons between PASD and other approaches can be found in Table III, demonstrating the impact of this text-based conditioning on image SR results. A similar concurrent work can be found with SeeSR [132]. XPSR [133] extends this idea by fusing different levels of semantic text encodings (high-level: the content of the image; low-level: the perception of overall quality, sharpness, noise level, and other distortions about the LR image).
@@
Figure 6: Comparison of the standard corruption space and I2SB. Instead of injecting noise to the clean image (initial state $\mathbf{z}_{0}$), the final state $\mathbf{z}_{T}$ is the degraded image.
V-E Corruption Space
Karras et al. [77] identified three pillars of DMs: the noise schedule, the network parameterization, and the sampling algorithm. Recently, many authors argued to consider also different types of corruption instead of pure Gaussian noise used during forward diffusion like Soft Score Matching [138], i.e., the starting point for backward diffusion or the target for the forward diffusion $\mathbf{z}_{T}$. Soft Score Matching directly incorporates the filtering process within the SGM, training the model to predict a clean image. Upon corruption, this predicted image aligns with the diffused observation. Note that $\mathbf{z}_{T}$ may be represented differently due to alternative state domains (e.g., latent, frequency, or residual). Cold Diffusion [139] presents another ingenious way of modifying the corruption space for DDPMs. It shows that the generative capability is not strongly dependent on the choice of image degradation. It reveals new experimental types of diffusion besides Gaussian noise can be effectively used, like animorphosis (i.e., human faces iteratively degrading to animal faces). The Image-to-Image Schrödinger Bridge (I2SB) goes in a similar direction but does not impose any assumptions on the underlying prior distributions [140]. In its diffusion process, the clean image represents the initial state, while the degraded image is the final state in both forward and backward diffusion processes. This is notable for its ability to provide a transparent and traceable path from a degraded image to its clean version, as illustrated in Figure 6. Consequently, it provides enhanced interpretability since the process between degraded and clean images is directly addressed, which is not commonly present in many DMs. Another benefit is its higher efficiency in backward diffusion since it requires fewer steps (often between 2 and 10) to achieve comparable performance. Its conditionality, however, limits its use specifically to paired data during training, which is unsuitable for unsupervised SR. While Cold Diffusion and I2SB show promising results for image restoration, an extensive and more detailed quantitative analysis of different corruption types for image SR remains an exciting and open research avenue. Another avenue for alternative corruption space is presented by Inversion by Direct Iteration (InDI) [141]. InDI delineates a direct mapping strategy, efficiently bridging the gap between the two quality spaces without the iterative refinement typically required by conventional diffusion processes. The intrinsic flexibility and the direct mapping capability of InDI propose intriguing possibilities for enhancing image quality, suggesting a potent avenue for research exploration. The potential integration of InDI’s principles with those of conditional DMs could offer substantial advancements in the field of image SR. A detailed examination and discussion of InDI within the broader scope of diffusion-based image enhancement could yield valuable insights and contribute significantly to the ongoing development of generative models in image processing.
V-F Color Shifting
V-G Architecture Designs for Denoising
The design of the denoising model in DMs offers a range of options. The majority of DMs adopt the use of U-Net, as noted in most literature [102]. SR3 [15], for instance, employs residual blocks from BigGAN [102] and re-scales skip connections by a factor of $\frac{1}{\sqrt{2}}$. SRDiff takes a similar approach [79], although it opts for vanilla residual blocks without the re-scaling of skip connections and uses a LR encoder to incorporate the information of the LR image during the backward diffusion. Whang et al. [117] exploit an initial predictor to combine the strengths of deterministic image SR models and DMs. It has the advantage that the DM only needs to learn the residuals that the deterministic image SR model (initial predictor) fails to predict, hence simplifying the learning target. Additionally, the removal of self-attention, positional encodings and group normalization from the SR3 U-Net enables their model to support arbitrary resolutions. An initial predictor is also employed in the wavelet-based approach DiWa [113]. Moreover, wavelet SR models, such as DWSR [112] – a simple sequence of convolution layers of depth 10 – are utilized for denoising prediction in the wavelet domain. In WaveDM [114], a deterministic U-Net predictor is used for the high-frequency band, while diffusion is applied in the low-frequency band.